I need help with this equation 2A/3 = 8 + 4A


Answer by stanbon(75887)

2A/3 = 8 + 4A
---------
Multiply thru by 3 to get:
2A = 24 + 12A
10A = -24
A = -12/5
=============
Cheers,
Stan H.